Please refer to illustration for question.
Find a QR factorization of the matrix A. 0 1 1 0 A = 1 1 -1 -1 1 1
Answer #1

Solution to Given : 4. 1 Rii Riz [2 a 9 R22 R23 where, 2, = a, a 1 R., = 119, 11 = 33 q = . Ru â 1/53 Ria = 9, 92 ** to do agSo, R₂3 Risa qa, = 0 q, ag - :-/3/11 Again å = Az - Rigg, - R234 o +3/11 O. + 2/1 1-2/11 Foto -4/11 7/11 So, R33 = 119 11 = √